# Wasserstein Distributionally Robust Shallow Convex Neural Networks

### v2

@@ 2.2 Shallow Convex Neural Networks
Let $\mathbf{X}\in\mathbb{R}^{N\times d}$ be a data matrix of $N$ feature vectors $\mathbf{x}_{j}\in\mathbb{R}^{d}\ \forall j\in\llbracket N\rrbracket$, and $\mathbf{y}\in\mathbb{R}^{N}$ be the corresponding one-dimensional labels. Define data samples as a feature-label pair: $\mathbf{z}_{j}=(\mathbf{x}_{j},y_{j})\ \forall j\in\llbracket N\rrbracket$. The non-convex training problem of a single-output feedforward shallow neural network (SNN) with $\operatorname{ReLU}$ activation functions is as follows:
@@
The convex reformulation of the SNN training problem (SNNT) is based on an enumeration process of all the possible $\operatorname{ReLU}$ activation patterns in the hidden layer for a fixed training dataset $\mathbf{X}$ (Pilanci and Ergen 2020). The enumeration is equivalent to generating all possible hyperplane arrangements passing through the origin and clustering the data with them. The set of all possible activation patterns a $\operatorname{ReLU}$ SNN can take for a training set $\mathbf{X}$ is given by:
@@
$\mathcal{D}_{\mathbf{X}}=\left\{\mathbf{D}=\operatorname{diag}(\text{1}(\mathbf{Xs}\succeq\mathbf{0})):\mathbf{s}\in\mathbb{R}^{d}\right\},$ (1)
@@
where $\mathbf{D}$ is a diagonal matrix with a possible activation pattern encoded on the diagonal.
@@
By introducing the cone-constrained variables $\bm{\nu},\bm{\omega}\in\mathbb{R}^{P\times d}$ and the diagonal matrices from $\mathcal{D}_{\mathbf{X}}$ilanci and Ergen (2020) have obtained a convex training procedure from which we can retrieve the previous non-convex weights under some conditions. The optimization problem is now defined as
@@
$p^{\star}=$ $\min_{\bm{\nu},\bm{\omega}}$ $\mathcal{L}\left(\sum_{i=1}^{P}\mathbf{D}_{i}\mathbf{X}(\bm{\nu}_{i}-\bm{\omega}_{i}),\mathbf{y}\right)$ (SCNNT)
@@
s.t. $\bm{\nu}_{i},\bm{\omega}_{i}\in\mathcal{K}_{i}\quad\forall i\in\llbracket P\rrbracket,$
@@
where $\mathcal{K}_{i}=\{\mathbf{u}\in\mathbb{R}^{d}:(2\mathbf{D}_{i}-\mathbf{I})\mathbf{X}\mathbf{u}\succeq\mathbf{0}\}$ and $P=\operatorname{card}(\mathcal{D}_{\mathbf{X}})$. For $m\geq m^{\star}=\sum_{i:\bm{\nu}_{i}\neq\mathbf{0}}^{P}1+\sum_{i:,\bm{\omega}_{i}\neq\mathbf{0}}^{P}1$, with $11\leq m^{\star}\leq N+$ 1, problems (SNNT) and (SCNNT) have identical optimal values (Pilanci and Ergen 2020). At least one optimal non-convex NN from (SNNT) can be assembled, using $m^{\star}$ neurons, with the optimal solution of (SCNNT) and the following mapping between the optimal weights of each problem:
[…]
@@
To make the problem computationally tractable, we can approximate the full hyperplane arrangements by sampling random vectors (Mishkin et al. 2022). For example, using a Gaussian sampling, we substitute $\mathcal{D}_{\mathbf{X}}$ by $\tilde{\mathcal{D}}_{\mathbf{X}}=\{\mathbf{D}_{i}=\operatorname{diag}(\text{1}(\mathbf{XS}_{i}\succeq\mathbf{0})):\mathbf{S}_{i}\thicksim\mathcal{N}_{d}(\bm{\mu},\bm{\Sigma}),\mathbf{D}_{i}\neq\mathbf{D}_{j\neq i},i\in\llbracket P\rrbracket\}$ where $\mathbf{S}_{i}$ is a random Gaussian vector of mean $\bm{\mu}\in\mathbb{R}^{d}$ and of covariance $\bm{\Sigma}\in\mathbb{R}^{d\times d}$, and such that $\mathbf{s}_{i}\in\mathbb{R}^{d}$ is its realization, $\forall i\in\llbracket P\rrbracket$. This approximation has the same minima as (SCNNT) if
@@
where in this case $\bm{\nu}^{\star},\bm{\omega}^{\star}$ are global optima of the sub-sampled convex problem and if the optimal activations are in the convex model, i.e., $\left\{\operatorname{diag}(\mathbf{XW}_{1i}^{\star}\geq 0):i\in\llbracket m\rrbracket\right\}\subseteq\tilde{\mathcal{D}}_{X}$ (Mishkin et al. 2022). Interested readers are referred to Pilanci and Ergen (2020) and Mishkin et al. (2022) for the detailed coverage of SCNNs.
We remark that if these conditions are met, for a feature sample $\mathbf{x}$ not included in the training dataset $\mathbf{X}$, then the non-convex formulation of the neural network gives the same output as the convex formulation when using the same set of realized sampling vectors $\mathcal{S}=\{\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{P}\}$. This notion will be useful in the next section.
If these conditions are not met, i.e., the SNN is not wide enough, we still obtain an approximately optimal network that can outperform conventional training methods as highlighted by Mishkin et al. (2022). By approximately optimal, we mean that it is optimized by a convex program that guarantees optimality while not having an exact mapping to the optimal non-convex neural network formulation. Thus, choosing the maximal width of the SCNN is a way to control the tractability of the convex training with respect to the available computational power and the size of the training dataset.
@@ 3 Wasserstein Distributionally Robust Shallow Convex Neural Networks
We make the following assumption on samples from $\mathbb{P}$.
@@
This is a common assumption in statistical machine learning which is also made by, e.g., Chen and Paschalidis (2020) and Mohajerin Esfahani and Kuhn (2018). In practice, in learning settings, we only have access to a finite amount of training samples from which we can indirectly observe $\mathbb{P}$. This motivates the use of a finite empirical distribution ^$\hat{\mathbb{P}}_{N}$ that approximates $\mathbb{P}$ (Kuhn et al. 2019). As such, we assume that our empirical samples are drawn independently from a consistent unknown distribution which is i.i.d. When formulating (TWDRO), only the independence of the samples is necessary to build ^$\hat{\mathbb{P}}_{N}$. The stronger identically distributed assumption is then only required to bound the out-of-sample performance, see Section 4. As a result, the later part of Assumption 1 implies that any distribution drift is neglected from the analysis. Recall (1) and let $d_{ji}=\mathbf{D}_{i}(j,j)=\text{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{i}\geq\mathbf{0}),\ \forall j\in\llbracket N\rrbracket,\ i\in\llbracket P\rrbracket$ such that $\mathbf{d}_{j}\in\{0,1\}^{P}$ is the vector encoding the activation state of $\mathbf{x}_{j}$ for each vector in $\mathcal{S}$. For the remainder of this section, we consider the modified features ${\hat{\mathbf{x}}}_{j}=\left(\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top})\right)\ \forall j\in\llbracket N\rrbracket$, which are a higher order basis expansion of the original features with their respective possible activation patterns.
@@
Suppose that the training loss function $\mathcal{L}$ is the $\ell_{1}$-norm, and let $\hat{\mathbf{z}}_{j}=\left(\hat{\mathbf{x}}_{j},y_{j}\right)\ j\in\llbracket N\rrbracket$ be modified samples. Then, the convex training procedure (SCNNT) is equivalent to a constrained linear regression problem.
@@
and is equivalent to a constrained linear regression problem with respect to the modified samples. Additionally, we note that the LASSO and Ridge regularized SCNN training problems are exactly equivalent to their respective linear regression formulation.
$p^{\star}_{\ell_{1},r}=$ $\min_{\bm{\nu},\bm{\omega}:\bm{\nu}_{i},\bm{\omega}_{i}\in\mathcal{K}_{i}\ \forall i\in\llbracket P\rrbracket}$ $\left\|\mathbf{y}-\sum_{D_{i}\in\mathcal{D}_{\mathbf{X}}}\mathbf{D}_{i}\mathbf{X}(\bm{\nu}_{i}-\bm{\omega}_{i})\right\|_{1}+\sum_{i=1}^{P}\left(r(\bm{\nu}_{i})+r(\bm{\omega}_{i})\right)$
@@
where $r(\cdot)=\|\cdot\|_{1}$ for LASSO regularization and $r(\cdot)=\|\cdot\|_{2}^{2}$ for Ridge regularization. $\square$
The $\ell_{1}$-norm loss function tends to be more robust against outliers than other regression objectives when training regression models (Chen and Paschalidis 2020). It is then a preferable choice in an application where distributional robustness is desired and will be used for the remainder of this work as it is also convenient in the WaDiRo reformulation. Nonetheless, we remark that Lemma 1 can be adapted to other separable loss functions, i.e., functions $f$ such that $f(\phi_{1}(x_{1})+\phi_{2}(x_{2})+\ldots\phi_{N}(x_{N}))=f(\phi_{1}(x_{1}))+f(\phi_{2}(x_{2}))+\ldots+f(\phi_{N}(x_{N}))$ where $x_{j}\ j\in\llbracket N\rrbracket$ are variables and $\phi_{j}\ j\in\llbracket N\rrbracket$ are arbitrary functions.
We remark that the transformed feature space preserves the independence of individual samples when conditioned on $\mathcal{S}=\{\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{P}\}$ because the mapping from the original feature space to the modified one depends only on individual samples and $\mathcal{S}$. Whether $\mathcal{S}$ is constructed by random sampling or a by deterministic hyperplane arrangement has no importance as long as these vectors are saved prior to training and are used throughout the training and evaluation. From this observation, we now introduce our second lemma.
@@
Suppose Assumption 1 holds, i.e., samples $\mathbf{z}\in\mathcal{Z}$ are i.i.d., then modified samples given the set $\mathcal{S}$, $\hat{\mathbf{z}}|\mathcal{S}\in\hat{\mathcal{Z}}$, are also i.i.d.
@@
where $\hat{\mathbf{x}}_{j}=f(\mathbf{x}_{j},\mathcal{S})$ is a basis expansion of the original feature space using the definition of the hyperplane generation (or approximation) and $g_{\bm{\beta}}(\hat{\mathbf{z}}):\hat{\mathcal{Z}}\times\mathbb{R}^{2Pd+1}\to\mathbb{R}$ is the modified objective function when applying $\bm{\beta}$ on a modified sample. We have showed that each $\hat{\mathbf{x}}_{j}$ depends solely on $\mathbf{x}_{j}$ and $\mathcal{S}$. Because $\mathcal{S}$ is obtained prior to training and the mapping $f$ is the same on every sample, $\hat{\mathbf{z}}|\mathcal{S}$ is also i.i.d.  $\square$
Following Lemmas 1 and 2, we consider the new empirical distribution of the modified samples conditioned on $\mathcal{S}$: $\hat{\mathbb{F}}_{N}^{\mathcal{S}}=\frac{1}{N}\sum_{j=1}^{N}\delta_{\hat{\mathbf{z}}_{j}|\mathcal{S}}(\hat{\mathbf{z}}|\mathcal{S})$, where $\hat{\mathbf{z}}_{j}=\left(\hat{\mathbf{x}}_{j},y_{j}\right)^{\top}$. As the number of modified samples grows, $\hat{\mathbb{F}}_{N}^{\mathcal{S}}$ converges to the true underlying distribution $\mathbb{F}^{\mathcal{S}}$.
To make our case for the similarity between samples of ^$\hat{\mathbb{P}}_{N}$ and $\hat{\mathbb{F}}_{N}^{\mathcal{S}}$, we provide a detailed example of the mapping between the samples of the two distributions.
@@
Let $\mathbf{z}_{j}=(\mathbf{x}_{j},y_{j})$ be any sample from ^$\hat{\mathbb{P}}_{N}$ and $\hat{\mathbf{z}}_{j}=(\hat{\mathbf{x}}_{j},y_{j})$ be the equivalent sample from $\hat{\mathbb{F}}_{N}^{\mathcal{S}}$. Suppose that $\operatorname{card}(\mathcal{S})=3$ and that $\mathbf{d}_{j}^{\top}=\begin{bmatrix}\text{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{1}\geq 0)&\text{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{2}\geq 0)&\text{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{3}\geq 0)\end{bmatrix}=\begin{bmatrix}1&0&1\end{bmatrix}$, then:
[…]

### v3

@@ 2.2 Shallow Convex Neural Networks
Let $\mathbf{X}\in\mathbb{R}^{N\times d}$ be a data matrix of $N$ feature vectors $\mathbf{x}_{j}\in\mathbb{R}^{d}\ \forall j\in\llbracket N\rrbracket$, and $\mathbf{y}\in\mathbb{R}^{N}$ be the corresponding one-dimensional labels. Define data samples as a feature-label pair: $\mathbf{z}_{j}=(\mathbf{x}_{j},y_{j})\ \forall j\in\llbracket N\rrbracket$. The non-convex training problem of a single-output feedforward shallow neural network (SNN) with $\operatorname{ReLU}$ activation functions, without regularization, and without bias weights is as follows:
@@
The convex reformulation of the SNN training problem (SNNT) is based on an enumeration process of all the possible $\operatorname{ReLU}$ activation patterns in the hidden layer for a fixed training dataset $\mathbf{X}$ (Pilanci and Ergen 2020). The enumeration is equivalent to generating all possible hyperplane arrangements passing through the origin and clustering the data with them. The set of all $\bar{P}$ possible activation patterns a $\operatorname{ReLU}$ SNN can take for a training set $\mathbf{X}$ is given by:
@@
$\mathcal{D}_{\mathbf{X}}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\{\mathbf{D}_{1},\mathbf{D}_{2},\ldots,\mathbf{D}_{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{P}}}}:\mathbf{D}_{i}\neq\mathbf{D}_{j\neq i},\ i,j\in\llbracket{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{P}}}\rrbracket\}\subset}}\left\{\mathbf{D}=\operatorname{diag}(\text{1}(\mathbf{Xs}\succeq\mathbf{0})):\mathbf{s}\in\mathbb{R}^{d}\right\},$ (1)
@@
where $\mathbf{D}$ is a diagonal matrix with a possible activation pattern encoded on the diagonal and $\mathbf{s}$ is the vector used to encode the activation pattern with respect to the dataset $\mathbf{X}$.
@@
By introducing the cone-constrained variables $\bm{\nu},\bm{\omega}\in\mathbb{R}^{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{P}}}\times d}$ and the diagonal matrices from $\mathcal{D}_{\mathbf{X}}$Pilanci and Ergen (2020) have obtained a convex training procedure from which we can retrieve the previous non-convex weights under some conditions. The optimization problem is now defined as
@@
$p^{\star}=$ $\min_{\bm{\nu},\bm{\omega}}$ $\mathcal{L}\left(\sum_{i=1}^{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{P}}}}\mathbf{D}_{i}\mathbf{X}(\bm{\nu}_{i}-\bm{\omega}_{i}),\mathbf{y}\right)$ (SCNNT)
@@
s.t. $\bm{\nu}_{i},\bm{\omega}_{i}\in\mathcal{K}_{i}\quad\forall i\in\llbracket{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{P}}}\rrbracket,$
@@
where $\mathcal{K}_{i}=\{\mathbf{u}\in\mathbb{R}^{d}:(2\mathbf{D}_{i}-\mathbf{I})\mathbf{X}\mathbf{u}\succeq\mathbf{0}\}$ and ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{P}}}=\operatorname{card}(\mathcal{D}_{\mathbf{X}})$. For $m\geq m^{\star}=\sum_{i:\bm{\nu}_{i}\neq\mathbf{0}}^{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{P}}}}1+\sum_{i:,\bm{\omega}_{i}\neq\mathbf{0}}^{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{P}}}}1$, with $1\leq m^{\star}\leq N+1$, problems (SNNT) and (SCNNT) have identical optimal values (Pilanci and Ergen 2020). At least one optimal non-convex NN from (SNNT) can be assembled, using $m^{\star}$ neurons, with the optimal solution of (SCNNT) and the following mapping between the optimal weights of each problem:
$\left(\mathbf{W}_{1i}^{\star},\mathbf{w}_{2i}^{\star}\right)=\left(\frac{\bm{\nu}_{i}^{\star}}{\sqrt{\left\|\bm{\nu}_{i}^{\star}\right\|_{2}}},\sqrt{\left\|\bm{\nu}_{i}^{\star}\right\|_{2}}\right),\quad$ $\text{ if }\quad\bm{\nu}_{i}^{\star}\neq\mathbf{0},\forall i\in\llbracket{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{P}}}\rrbracket$ (2)
$\left(\mathbf{W}_{1(i+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{P}}})}^{\star},\mathbf{w}_{2(i+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{P}}})}^{\star}\right)=\left(\frac{\bm{\omega}_{i}^{\star}}{\sqrt{\left\|\bm{\omega}_{i}^{\star}\right\|_{2}}},-\sqrt{\left\|\bm{\omega}_{i}^{\star}\right\|_{2}}\right),\quad$ $\text{ if }\quad\bm{\omega}_{i}^{\star}\neq\mathbf{0},\forall i\in\llbracket{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{P}}}\rrbracket.$ (3)
@@
To make the problem computationally tractable, we can approximate the full hyperplane arrangements by sampling random vectors (Mishkin et al. 2022). For example, using a Gaussian sampling, we approximate $\mathcal{D}_{\mathbf{X}}$ by $\tilde{\mathcal{D}}_{\mathbf{X}}=\{\mathbf{D}_{i}=\operatorname{diag}(\text{1}(\mathbf{XS}\succeq\mathbf{0})):\mathbf{S}\thicksim\mathcal{N}_{d}(\bm{\mu},\bm{\Sigma}),\mathbf{D}_{i}\neq\mathbf{D}_{j\neq i},\ i,j\in\llbracket P\rrbracket\}$ where $P\leq\bar{P}$, $\mathbf{S}$ is a random Gaussian vector of mean $\bm{\mu}\in\mathbb{R}^{d}$ and of covariance $\bm{\Sigma}\in\mathbb{R}^{d\times d}$, and such that $\mathbf{s}_{i}\in\mathbb{R}^{d}$ is the realization generating matrix $\mathbf{D}_{i}$, $\forall i\in\llbracket P\rrbracket$. We note that other sampling methods can be used, e.g., uniform sampling, and they may possibly lead to richer sets $\tilde{\mathcal{D}}_{\mathbf{X}}$. This approximation has the same minima as (SCNNT) if
@@
where in this case $\bm{\nu}^{\star},\bm{\omega}^{\star}$ are global optima of the sub-sampled convex problem and if the optimal activations are in the convex model, i.e., $\left\{\operatorname{diag}(\mathbf{XW}_{1i}^{\star}\geq 0):i\in\llbracket m\rrbracket\right\}\subseteq\tilde{\mathcal{D}}_{\mathbf{X}}$ (Mishkin et al. 2022). Interested readers are referred to Pilanci and Ergen (2020) and Mishkin et al. (2022) for the detailed coverage of SCNNs.
We remark that if these conditions are met, for a feature sample $\mathbf{x}$ not included in the training dataset $\mathbf{X}$, then the non-convex formulation of the neural network gives the same output as the convex formulation when using the same set of realized sampling vectors $\mathcal{S}=\{\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{P}\}$. We also observe empirically that as the cardinality of $\tilde{\mathcal{D}}_{\mathbf{X}}$ grows, both formulation converges to the same outputs for a given input. This notion will be useful in the next section.
If these conditions are not met, i.e., the SNN is not wide enough, we still obtain an approximately optimal network that can outperform conventional training methods as highlighted by Mishkin et al. (2022). By approximate optimality, we mean that it is optimized by a convex program that guarantees optimality while not having an exact mapping to the optimal non-convex neural network formulation. Thus, imposing the maximal width of the SCNN, which depends on the cardinality of $\tilde{\mathcal{D}}_{\mathbf{X}}$, is a way to control the tractability of the convex training with respect to the available computational power and the size of the training dataset. Because generating the full hyperplane arrangement is impractical in most settings, the notation of the subsequent sections is aligned with the use of $\tilde{\mathcal{D}}_{\mathbf{X}}$.
@@ 3 Wasserstein Distributionally Robust Shallow Convex Neural Networks
We make the following assumption on samples from P.
@@
This is a common assumption in statistical machine learning which is also made by, e.g., Chen and Paschalidis (2020) and Mohajerin Esfahani and Kuhn (2018). In practice, in learning settings, we only have access to a finite amount of training samples from which we can indirectly observe P. This motivates the use of a finite empirical distribution $\hat{\text{P}}_{N}$ that approximates P (Kuhn et al. 2019). As such, we assume that our empirical samples are drawn independently from a consistent unknown distribution with i.i.d. samples. When formulating (TWDRO), only the independence of the samples is necessary to build $\hat{\text{P}}_{N}$. The stronger identically distributed assumption is then only required to bound the out-of-sample performance, see Section 4. As a result, the latter part of Assumption 1 implies that any distribution drift is neglected from the analysis. While Assumption 1 is not always preserved in practice, e.g., time-series forecasting, it is required for the theoretical analysis. Recall (1) and let $d_{ji}=\mathbf{D}_{i}(j,j)=\text{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{i}\geq\mathbf{0}),\ \forall j\in\llbracket N\rrbracket,\ i\in\llbracket P\rrbracket$ such that $\mathbf{d}_{j}\in\{0,1\}^{P}$ is the vector encoding the activation state of $\mathbf{x}_{j}$ for each vector in $\mathcal{S}$. For the remainder of this section, we consider the modified features ${\hat{\mathbf{x}}}_{j}=\left(\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top})\right)\ \forall j\in\llbracket N\rrbracket$, which are a higher order basis expansion of the original features with their respective possible activation patterns. This basis expansion is justified in the proof of the following lemma.
@@
Suppose that the training loss function $\mathcal{L}$ is the $\ell_{1}$-norm, and let $\hat{\mathbf{z}}_{j}=\left(\hat{\mathbf{x}}_{j},y_{j}\right)\in\hat{\mathcal{Z}},\ j\in\llbracket N\rrbracket$ be modified samples. Then, the convex training procedure (SCNNT) is equivalent to a constrained linear regression problem.
@@
and is equivalent to a constrained linear regression problem with respect to the modified samples. $\square$
Additionally, we note that the LASSO and ridge-regularized SCNN training problems are exactly equivalent to their respective linear regression formulation:
$p^{\star}_{\ell_{1},r}=$ $\min_{\bm{\nu},\bm{\omega}:\bm{\nu}_{i},\bm{\omega}_{i}\in\mathcal{K}_{i}\ \forall i\in\llbracket P\rrbracket}$ $\left\|\mathbf{y}-\sum_{i=1}^{P}\mathbf{D}_{i}\mathbf{X}(\bm{\nu}_{i}-\bm{\omega}_{i})\right\|_{1}+\sum_{i=1}^{P}\left(r(\bm{\nu}_{i})+r(\bm{\omega}_{i})\right)$
@@
where $r(\cdot)=\|\cdot\|_{1}$ for LASSO regularization and $r(\cdot)=\|\cdot\|_{2}^{2}$ for ridge regularization.
The $\ell_{1}$-norm loss function tends to be more robust against outliers than other regression objectives when training regression models (Chen and Paschalidis 2020). It is then a preferable choice in an application where distributional robustness is desired and will be used for the remainder of this work as it is also convenient in the WaDiRo reformulation. Nonetheless, we remark that Lemma 1 can be adapted to other separable loss functions, i.e., functions $f$ such that $f(\phi_{1}(x_{1})+\phi_{2}(x_{2})+\ldots\phi_{N}(x_{N}))=f(\phi_{1}(x_{1}))+f(\phi_{2}(x_{2}))+\ldots+f(\phi_{N}(x_{N}))$ where $x_{j}\ j\in\llbracket N\rrbracket$ are variables and $\phi_{j}\ j\in\llbracket N\rrbracket$ are arbitrary functions. Other commonly used separable loss functions include the $\ell_{p}$-norm to the power of $p$, i.e., $\|\cdot\|_{p}^{p}$, for $p\in[1,+\infty)$.
We remark that the transformed feature space preserves the independence of individual samples when conditioned on $\mathcal{S}=\{\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{P}\}$ because the mapping from the original feature space to the modified one depends only on individual samples and $\mathcal{S}$. Whether $\mathcal{S}$ is constructed by random sampling or by a deterministic hyperplane arrangement has no importance as long as these vectors are saved prior to training and are used throughout the training and evaluation. From this observation, we now introduce our second lemma.
@@
Suppose Assumption 1 holds, i.e., samples $\mathbf{z}=(\mathbf{x},y)\in\mathcal{Z}$ are i.i.d., then modified samples $\hat{\mathbf{z}}=(\hat{\mathbf{x}},y)$ given the set $\mathcal{S}$, $\hat{\mathbf{z}}|\mathcal{S}\in\hat{\mathcal{Z}}$, are also i.i.d.
@@
where $\hat{\mathbf{x}}_{j}=f(\mathbf{x}_{j},\mathcal{S})$ is a basis expansion of the original feature space using the definition of the hyperplane generation (or approximation) and $g_{\bm{\beta}}(\hat{\mathbf{z}}):\hat{\mathcal{Z}}\times\mathbb{R}^{2Pd+1}\to\mathbb{R}$ is the modified objective function when applying $\bm{\beta}$ on a modified sample. We have showed that each $\hat{\mathbf{x}}_{j}$ depends solely on $\mathbf{x}_{j}$ and $\mathcal{S}$. Because $\mathcal{S}$ is obtained before training, either by random sampling or by a deterministic hyperplane arrangement which depends on the whole training set $\mathbf{X}$, and the mapping $f$ is the same on every sample, $\hat{\mathbf{z}}|\mathcal{S}$ is also i.i.d.  $\square$
Following Lemmas 1 and 2, we consider the new empirical distribution of the modified samples conditioned on $\mathcal{S}$: $\hat{\text{F}}_{N}^{\mathcal{S}}=\frac{1}{N}\sum_{j=1}^{N}\delta_{\hat{\mathbf{z}}_{j}|\mathcal{S}}$, where $\hat{\mathbf{z}}_{j}=\left(\hat{\mathbf{x}}_{j},y_{j}\right)^{\top}$. As the number of modified samples grows, $\hat{\text{F}}_{N}^{\mathcal{S}}$ converges to the true underlying distribution $\text{F}^{\mathcal{S}}$.
To make our case for the similarity between samples of $\hat{\text{P}}_{N}$ and $\hat{\text{F}}_{N}^{\mathcal{S}}$, we provide a detailed example of the mapping between the samples of the two distributions.
@@
Let $\mathbf{z}_{j}=(\mathbf{x}_{j},y_{j})$ be any sample from $\hat{\text{P}}_{N}$ and $\hat{\mathbf{z}}_{j}=(\hat{\mathbf{x}}_{j},y_{j})$ be the equivalent sample from $\hat{\text{F}}_{N}^{\mathcal{S}}$. Suppose that $\operatorname{card}(\mathcal{S})=3$ and that $\mathbf{d}_{j}^{\top}=\begin{bmatrix}\text{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{1}\geq 0)&\text{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{2}\geq 0)&\text{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{3}\geq 0)\end{bmatrix}=\begin{bmatrix}1&0&1\end{bmatrix}$, then:
$\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top})$ $=\operatorname{vec}\left(\mathbf{x}_{j}\begin{bmatrix}1&0&1\end{bmatrix}\right)=\operatorname{vec}\left(\begin{bmatrix}\mathbf{x}_{j}&\mathbf{0}_{d}&\mathbf{x}_{j}\end{bmatrix}\right)=\operatorname{vec}\left(\begin{bmatrix}{x}_{j1}&0&{x}_{j1}\\ {x}_{j2}&0&{x}_{j2}\\ \vdots&\vdots&\vdots\\ {x}_{jd}&0&{x}_{jd}\\ \end{bmatrix}\right)=(\mathbf{x}_{j},\ \mathbf{0}_{d},\ \mathbf{x}_{j}),$